# Trace Reconstruction of First-Order Reed-Muller Codewords Using Run Statistics

### v1

@@ V Proof of Lemma 3
0000111100001111 $\frac{8929}{1024}$ $\frac{1811}{8192}$ $\frac{58275}{32768}$
@@
$0011$ $\frac{23593}{4096}$ $\frac{655}{1024}$ $\frac{44559}{32768}$
@@
0011110000111100 $\frac{251033}{32768}$ $\frac{41939}{32768}$ $\frac{5895}{8192}$
@@
$0101$ $\frac{36409}{8192}$ $\frac{7279}{8192}$ $\frac{36403}{32768}$
@@
0101101001011010 $\frac{164861}{32768}$ $\frac{35591}{32768}$ $\frac{3741}{4096}$
@@
$0110$ $\frac{175637}{32768}$ $\frac{34067}{32768}$ $\frac{7863}{8192}$
@@
0110100101101001 $\frac{42179}{8192}$ $\frac{3967}{4096}$ $\frac{33783}{32768}$
@@
[…]
@@
[…]
@@
[…]
@@
$|\alpha_{x}-\delta_{x}|$ $\leq\ \alpha_{x}+\delta_{x}\ =\ a(n)\ \leq\ n$ (6)
[…]
@@
Thus, (C4) holds for all codewords in RM($m+1,1$). ∎
@@ VI Conclusion
In this paper, we derived, for a given binary sequence $x$, an expression for the expected number of runs, $\mathbb{E}[R_{x}]$, in a random trace obtained by passing $x$ through an i.i.d. deletion channel with deletion probability $q$. We then proved that, when $q=\frac{1}{2}$, the average number of runs in $\widetilde{O}(n)$ traces can be used as a statistic to distinguish (w.h.p.) between the codewords of a first-order Reed-Muller code of blocklength $n$. To prove this, we used a key property of $\mathrm{RM}(m,1)$: for distinct codewords $x,y\in\mathrm{RM}(m,1)$ that share the same first bit, $|\mathbb{E}[R_{x}]-\mathbb{E}[R_{y}]|\geq\Delta$, for some constant $\Delta>0$ that does not depend on $m$. This property does not hold, for example, for higher-order RM codes $\mathrm{RM}(m,r)$, $r\geq 2$, so $\widetilde{O}(n)$ traces will not be sufficient to reconstruct $x\in\mathrm{RM}(m,r)$ (up to complementation) solely using the average number of runs in the traces as a statistic. Nonetheless, it should be possible to refine this statistic further or combine it with other statistics to better effect.
@@ Appendix B Proofs of (C1)–(C3) in Cases 2–4
[…]
@@ B-A Proof of (C1)–(C3) in Case 2
$|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|$ $=\textstyle{\left|\bigl{(}2+(\frac{1}{2})^{n}\bigr{)}|\alpha_{x}-\alpha_{y}|+\beta_{x}-\beta_{y}\right|}$
$|\beta_{\widehat{x}}-\beta_{\widehat{y}}|$ $=\textstyle{\left|\bigl{(}1+(\frac{1}{2})^{n-1}\bigr{)}(\beta_{x}-\beta_{y})+(\frac{1}{2})^{n}(\alpha_{x}-\alpha_{y})\right|}$
@@
    $|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|$ $=\textstyle{\bigl{(}2+(\frac{1}{2})^{n}\bigr{)}|\alpha_{x}-\alpha_{y}|+|\beta_{x}-\beta_{y}|}$ (16)
@@
    $|\beta_{\widehat{x}}-\beta_{\widehat{y}}|$ $=\textstyle{\bigl{(}1+(\frac{1}{2})^{n-1}\bigr{)}|\beta_{x}-\beta_{y}|+{(\frac{1}{2})}^{n}|\alpha_{x}-\alpha_{y}|}$
@@
    $\stackrel{{\scriptstyle\mathrm{(v)}}}{{\leq}}|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|$
    where we get the inequality labeled (iv) using the induction hypothesis that (C2) holds for RM$(m,1)$, whereas we obtain the inequality labeled (v) from (16).
@@
    $|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|$ $\geq\textstyle{\bigl{(}1+(\frac{1}{2})^{n}\bigr{)}|\alpha_{x}-\alpha_{y}|}$ (17)
    where we get the above inequality using the induction hypothesis that (C2) holds for RM$(m,1)$.
@@
    $\ |\beta_{\widehat{x}}-\beta_{\widehat{y}}|$ $\leq\textstyle{|\beta_{x}\!-\beta_{y}|+(\frac{1}{2})^{n}\left|(\alpha_{x}\!-\alpha_{y})+2(\beta_{x}\!-\beta_{y})\right|}$
@@
    $\stackrel{{\scriptstyle\mathrm{(vii)}}}{{\leq}}|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|$
    where the inequality labeled (vi) is obtained using the induction hypothesis that (C2) holds for RM$(m,1)$, whereas we get the inequality labeled (vii) from (17).
@@
$|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|$ $\geq|\alpha_{x}-\alpha_{y}|$ (18)
@@
where the inequality labeled (viii) is obtained using the induction hypothesis that (C1) holds for RM$(m,1)$.
@@
$\stackrel{{\scriptstyle\mathrm{(ix)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\cdot k+(\frac{1}{2})^{n}\cdot 3n}$
$\stackrel{{\scriptstyle\mathrm{(x)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|-0.06+4\sum\limits_{k=16}^{n}(\frac{1}{2})^{k}\cdot k}$
where we get the inequality labeled (ix) using (10), (12) and the induction hypothesis that (C3) holds for RM$(m,1)$. From (18), we obtain the inequality labeled (x).
@@ B-B Proof of (C1)–(C3) in Case 3
=$\stackrel{{\scriptstyle\mathrm{(xi)}}}{{=}}\textstyle|2(\alpha_{x}-\alpha_{y})-(\frac{1}{2})^{n}(\alpha_{x}-\alpha_{y})-(\beta_{x}-\beta_{y})|$
@@
$\stackrel{{\scriptstyle\mathrm{(xiii)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\cdot k-(\frac{1}{2})^{n}\cdot n}$
$\geq\textstyle{0.06-3\sum\limits_{k=16}^{n}(\frac{1}{2})^{k}\cdot k}$
where we get the inequality labeled (xi) using (13) and (14). We get the inequality labeled (xii) using the induction hypothesis that (C2) holds for RM$(m,1)$. We get the inequality labeled (xiii) using (9) and the induction hypothesis that (C1) holds for RM$(m,1)$.
@@
=$\stackrel{{\scriptstyle\mathrm{(xiv)}}}{{=}}\textstyle{|\beta_{x}\!-\beta_{y}\!-(\frac{1}{2})^{n-1}(\beta_{x}\!-\beta_{y})+(\frac{1}{2})^{n}(\alpha_{x}\!-\alpha_{y})|}$
$\leq\textstyle{\left(1-(\frac{1}{2})^{n-1}\right)\cdot|\beta_{x}-\beta_{y}|+(\frac{1}{2})^{n}\cdot|\alpha_{x}-\alpha_{y}|}$
$\stackrel{{\scriptstyle\mathrm{(xv)}}}{{\leq}}\textstyle{\bigl{(}1-(\frac{1}{2})^{n}\bigr{)}\cdot|\alpha_{x}-\alpha_{y}|}$
@@
where we get the inequality labeled (xiv) using (13) and (14). The inequality labeled (xv) is obtained using the induction hypothesis that (C2) holds for RM$(m,1)$, whereas we get the inequality labeled (xvi) from (19).
@@
$=\textstyle{|\gamma_{x}-\beta_{y}+(\frac{1}{2})^{n-1}(\beta_{x}-\gamma_{y})+(\frac{1}{2})^{n}(\delta_{x}-\alpha_{y})-(\frac{1}{2})^{n}\cdot n|}$
$\stackrel{{\scriptstyle\mathrm{(xvii)}}}{{\leq}}\textstyle{|\beta_{x}-\gamma_{y}|+(\frac{1}{2})^{n}\cdot 3n}$
$\stackrel{{\scriptstyle\mathrm{(xviii)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\!\cdot k+(\frac{1}{2})^{n}\!\cdot 3n}$
$\stackrel{{\scriptstyle\mathrm{(xix)}}}{{\leq}}\textstyle{|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|+(\frac{1}{2})^{n}\cdot n-0.06+4\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\cdot k+(\frac{1}{2})^{n}\cdot 3n}$
$\leq\textstyle{|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n}(\frac{1}{2})^{k}\cdot k}$
where we get the inequality labeled (xvii) from (10), (12) and (15). The inequality labeled (xviii) is obtained using the induction hypothesis that (C3) holds for RM$(m,1)$. From (19), the inequality labeled (xix) is obtained.
@@ B-C Proof of (C1)–(C3) in Case 4
$\geq\textstyle{2|\alpha_{x}-\alpha_{y}|-|\beta_{x}-\gamma_{y}|-(\frac{1}{2})^{n}(\alpha_{x}-\delta_{y})-n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xx)}}}{{\geq}}\textstyle{|\alpha_{x}\!-\alpha_{y}|+0.06-4\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\!\cdot k-2n(\frac{1}{2})^{n}}$ (20)
$\stackrel{{\scriptstyle\mathrm{(xxi)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\!\cdot k+0.06-4\sum\limits_{k=16}^{n}(\frac{1}{2})^{k}\!\cdot k}$
@@
where to obtain the inequality labeled (xx), we used (10) and the induction hypothesis that (C3) holds for RM$(m,1)$. The inequality labeled (xxi) is obtained using the induction hypothesis that (C1) holds for RM$(m,1)$. The inequality (xxii) is based on the fact that $\sum\limits_{k=16}^{n}k\,(\frac{1}{2})^{k}\leq\sum\limits_{k=16}^{\infty}k\,(\frac{1}{2})^{k}=17\cdot(\frac{1}{2})^{15}<0.001$.
@@
$\stackrel{{\scriptstyle\mathrm{(xxiii)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\!\!\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\!\cdot k+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xxiv)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.12+8\!\!\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\!\cdot k+5n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.12+8\sum\limits_{k=16}^{n}(\frac{1}{2})^{k}\cdot k}$
@@
where to obtain the inequality labeled (xxiii), we used (10), (11) and the induction hypothesis that (C3) holds for RM$(m,1)$. The inequality labeled (xxiv) follows directly from (20), whereas we get the inequality labeled (xv) based on the fact that $\sum\limits_{k=16}^{n}k\,(\frac{1}{2})^{k}\leq 0.001$.
@@
$\stackrel{{\scriptstyle\mathrm{(xxvii)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n/2}(\frac{1}{2})^{k}\!\cdot k+2n(\frac{1}{2})^{n}+2n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n}(\frac{1}{2})^{k}\cdot k}$
where to obtain the inequality labeled (xxvi), we used (9), (12) and the induction hypothesis that (C2) holds for RM$(m,1)$. From (20), we get the inequality labeled (xxvii).

### v3

@@ V Proof of Lemma 3
$00001111$ $\frac{8929}{1024}$ $\frac{1811}{8192}$ $\frac{58275}{32768}$
@@
$00110011$ $\frac{23593}{4096}$ $\frac{655}{1024}$ $\frac{44559}{32768}$
@@
$00111100$ $\frac{251033}{32768}$ $\frac{41939}{32768}$ $\frac{5895}{8192}$
@@
$01010101$ $\frac{36409}{8192}$ $\frac{7279}{8192}$ $\frac{36403}{32768}$
@@
$01011010$ $\frac{164861}{32768}$ $\frac{35591}{32768}$ $\frac{3741}{4096}$
@@
$01100110$ $\frac{175637}{32768}$ $\frac{34067}{32768}$ $\frac{7863}{8192}$
@@
$01101001$ $\frac{42179}{8192}$ $\frac{3967}{4096}$ $\frac{33783}{32768}$
@@
    $\alpha_{\widehat{x}}\ =\ 2\alpha_{x}+\beta_{x}+{(\frac{1}{2})}^{n}\alpha_{x}+n{(\frac{1}{2})}^{n}$
@@
    $\alpha_{\widecheck{x}}\ =\ 2\alpha_{x}+\gamma_{x}+{(\frac{1}{2})}^{n}\delta_{x}$
@@
$b(n)$ $=\sum_{k=1}^{n-1}k{\biggl(\frac{1}{2}\biggr)}^{k}=\ 2-(n+1){\biggl(\frac{1}{2}\biggr)}^{n-1}\ \leq\ 2$ (5)
@@
$|\alpha_{x}-\delta_{x}|\$$\leq\ \alpha_{x}+\delta_{x}\ =\ a(n)\ \leq\ n$ (6)
$|\beta_{x}-\gamma_{x}|\$$\leq\ \beta_{x}+\gamma_{x}\ =\ b(n)\ \leq\ 2.$ (7)
@@
Thus, (C4) holds for all codewords in $\mathrm{RM}(m+1,1)$. ∎
@@ VI Conclusion
In this paper, we derived, for a given binary sequence $x$, an expression for the expected number of runs, $\mathbb{E}[R_{x}]$, in a random trace obtained by passing $x$ through an i.i.d. deletion channel with deletion probability $q$. We then proved that, when $q=\frac{1}{2}$, the average number of runs in $\widetilde{O}(n^{2})$ traces can be used as a statistic to distinguish (w.h.p.) between the codewords of a first-order Reed-Muller code of blocklength $n$. To prove this, we used a key property of $\mathrm{RM}(m,1)$: for distinct codewords $x,y\in\mathrm{RM}(m,1)$ that share the same first bit, $|\mathbb{E}[R_{x}]-\mathbb{E}[R_{y}]|\geq\Delta$, for some constant $\Delta>0$ that does not depend on $m$. This property does not hold, for example, for higher-order RM codes $\mathrm{RM}(m,r)$, $r\geq 2$, so $\widetilde{O}(n^{2})$ traces might not be sufficient to reconstruct $x\in\mathrm{RM}(m,r)$ (up to complementation) solely using the average number of runs in the traces as a statistic. Nonetheless, it should be possible to refine this statistic further or combine it with other statistics to better effect.
@@ Appendix B Proofs of (C1)–(C3) in Cases 2–4
First let us demonstrate a few more properties of coefficients $\alpha,\beta,\gamma,\delta$ that we will be using in our proofs. As can be observed from their definitions, coefficients $\alpha,\beta,\gamma,\delta$ are non-negative. Thus, for a sequence $x$ of length $n$, using (4) and (5), we get $0\leq\alpha_{x}\leq n,\ 0\leq\delta_{x}\leq n,\ 0\leq\beta_{x}\leq 2,\ 0\leq\gamma_{x}\leq 2$. As a result, for any two sequences $x$ and $y$ of length $n$,
@@ B-A Proof of (C1)–(C3) in Case 2
[…]
@@
[…]
@@
[…]
@@
[…]
    where we derive inequality labeled (iv) by applying the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$, while inequality labeled (v) is obtained directly from (16).
@@
[…]
    where we get the above inequality using the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$.
@@
[…]
@@
[…]
    where the inequality labeled (vi) is derived using the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$, while the inequality labeled (vii) follows directly from (17).
@@
[…]
@@
where the inequality labeled (viii) is obtained using the induction hypothesis that (C1) holds for $\mathrm{RM}(m,1)$.
@@
$\stackrel{{\scriptstyle\mathrm{(ix)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(x)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|-0.06+4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
Using (10), (12), and the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$, we establish the inequality labeled (ix). Meanwhile, (18) leads to the inequality labeled (x).
@@ B-B Proof of (C1)–(C3) in Case 3
$\stackrel{{\scriptstyle\mathrm{(xi)}}}{{=}}\textstyle|2(\alpha_{x}-\alpha_{y})-(\frac{1}{2})^{n}(\alpha_{x}-\alpha_{y})-(\beta_{x}-\beta_{y})|$
@@
$\stackrel{{\scriptstyle\mathrm{(xiii)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}-n(\frac{1}{2})^{n}}$
$\geq\textstyle{0.06-3\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where the equality labeled (xi) is derived from (13) and (14). The inequality labeled (xii) follows from the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$. Finally, the inequality labeled (xiii) is obtained using (9) and the induction hypothesis that (C1) holds for $\mathrm{RM}(m,1)$.
@@
$\stackrel{{\scriptstyle\mathrm{(xiv)}}}{{=}}\textstyle{|\beta_{x}\!-\!\beta_{y}\!-(\frac{1}{2})^{n-1}(\beta_{x}\!-\!\beta_{y})\!+(\frac{1}{2})^{n}(\alpha_{x}\!-\!\alpha_{y})|}$
$\leq\textstyle{\left(1-(\frac{1}{2})^{n-1}\right)|\beta_{x}-\beta_{y}|+(\frac{1}{2})^{n}|\alpha_{x}-\alpha_{y}|}$
$\stackrel{{\scriptstyle\mathrm{(xv)}}}{{\leq}}\textstyle{\bigl(1-(\frac{1}{2})^{n}\bigr)|\alpha_{x}-\alpha_{y}|}$
@@
where the equality (xiv) is obtained directly from (13) and (14). The inequality labeled (xv) is established using the induction hypothesis that (C2) applies to $\mathrm{RM}(m,1)$, while inequality (xvi) results from (19).
@@
$=\textstyle{|\gamma_{x}-\beta_{y}+(\frac{1}{2})^{n-1}(\beta_{x}-\gamma_{y})+(\frac{1}{2})^{n}(\delta_{x}-\alpha_{y})-n(\frac{1}{2})^{n}|}$
$\stackrel{{\scriptstyle\mathrm{(xvii)}}}{{\leq}}\textstyle{|\beta_{x}-\gamma_{y}|+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xviii)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xix)}}}{{\leq}}\textstyle{|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|+n(\frac{1}{2})^{n}-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where the inequality (xvii) is obtained using (10), (12), and (15). The inequality labeled (xviii) is established by applying the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$, while inequality (xix) results from (9) and (19).
@@ B-C Proof of (C1)–(C3) in Case 4
$\geq\textstyle{2|\alpha_{x}-\alpha_{y}|-|\beta_{x}-\gamma_{y}|-(\frac{1}{2})^{n}|\alpha_{x}-\delta_{y}|-n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xx)}}}{{\geq}}\textstyle{|\alpha_{x}\!-\alpha_{y}|+0.06-4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}-2n(\frac{1}{2})^{n}}$ (20)
$\stackrel{{\scriptstyle\mathrm{(xxi)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+0.06-4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
@@
where the inequality labeled (xx) is derived using (10) along with the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$. The inequality labeled (xxi) is established based on the induction hypothesis that (C1) holds for $\mathrm{RM}(m,1)$. Finally, inequality (xxii) follows from the fact that $\sum\limits_{k=16}^{n}k\,(\frac{1}{2})^{k}\leq\sum\limits_{k=16}^{\infty}k\,(\frac{1}{2})^{k}=17\cdot(\frac{1}{2})^{15}<0.001$.
@@
$\stackrel{{\scriptstyle\mathrm{(xxiii)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xxiv)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.12+8\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+5n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.12+8\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
@@
where to obtain the inequality labeled (xxiii), we used (10), (11) and the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$. The inequality labeled (xxiv) follows directly from (20), whereas we get the inequality labeled (xxv) based on the fact that $\sum\limits_{k=16}^{n}k\,(\frac{1}{2})^{k}\leq 0.001$.
@@
$\stackrel{{\scriptstyle\mathrm{(xxvii)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+2n(\frac{1}{2})^{n}+2n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where to obtain the inequality labeled (xxvi), we used (9), (12) and the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$. Meanwhile, the inequality labeled (xxvii) follows directly from (20).